OpenStudy (dottiecat):

What is the equation of a line, in point-slope form, that passes through (-2, -6) and has a slope of 1/3 A y+2=1/3(x+6) B y+6=1/3(x+2) C y−6=1/3(x−2) D y−2=1/3(x−6)

OpenStudy (dottiecat):

I'm having trouble trying to find the answer

OpenStudy (mathstudent55):

Start with the point-slope form of the equation of a line. Can you show it?

OpenStudy (dottiecat):

I do not understand. Sorry...

OpenStudy (dottiecat):

oh wait,

OpenStudy (mathstudent55):

Ok. Let's cover the basics. The point-slope form of the equation of a line is \(y - y_1 = m(x - x_1)\) where \((x_1, y_1)\) is a point on the line and \(m\) is the slope of the line.

OpenStudy (mathstudent55):

Ok so far?

OpenStudy (dottiecat):

Yes, I'm used to calling it slope formula.

OpenStudy (dottiecat):

Just a second, I might've figured out the answer.

OpenStudy (mathstudent55):

The slope formula is a different thing. That is the formula you use to find the slope of a line given two points on the line. Back to this problem. You are given a point. You have point (-2, -6). The point in the equation is \((x_1, y_1)\) That means \(x_1 = -2\), and \(y_1 = -6\) You are also given the slope \(\dfrac{1}{3} \). Just plug those values into the point-slope form I gave you above, and that is the equation of the line.

OpenStudy (mathstudent55):